OpenStudy (anonymous):

Find the x and y intercepts and graph the equation plotting the intercepts -3x-7y=21

OpenStudy (saifoo.khan):

When you have to find the x-intercept, make y = 0 When you have to find the y-intercept, make x = 0

OpenStudy (anonymous):

did you understand @ronholloway

OpenStudy (anonymous):

no I am confused

OpenStudy (anonymous):

okay as saifoo said any line passing through the x axis has coordinates (x,0) and any line passing through the y axis has coordinates (0,y)

OpenStudy (anonymous):

so the points should satisfy the equation of the line.... putting the value of the first pt.(x,0) -3x-7*0=21 -3x=21 x=-7 which is the x-intercept.... putting the value of the second pt.(0,y) -3*0-7*y=21 -7y=21 y=-3 which is the y-intercept....

OpenStudy (anonymous):

ok I we did get the same answer for the y-intercept; i think i have it now

OpenStudy (anonymous):

good...

OpenStudy (anonymous):

thanks